# Charges, complex structures, and perturbations of instantons

### v2

@@ 2.2. The self-dual Weyl spinor
In section 5 we shall be interested in eigenspinors for $\Psi_{ABCD}$. A detailed study of this can be found in [40, §8.3]. Here we collect some facts adapted to Riemannian signature, for later use in section 5.
@@
where the $\phi^{i}_{AB}$’s are orthonormal, $I^{ABCD}\phi^{i}_{AB}\phi^{j}_{CD}=\delta^{ij}$. From [40, Eq. (8.3.8), (8.3.11)], the eigenvalues satisfy the following identities:
@@ 3.1. Charges for Hermitian instantons
The proof of (3.3) is straightforward if one uses a ‘conformal-GHP connection’ [6]: this is an operator $\mathcal{C}_{AA^{\prime}}$ which generalizes the ordinary GHP connection [39, §4.12] to be invariant under combined conformal transformations and rescalings of the spinors $o_{A},\iota_{A}$, by taking into account both the conformal and the GHP weights of the objects it acts on. The necessary background is given in appendix A. Since $\zeta_{AB}$ has conformal weight $w=1$, then $\zeta^{AB}$ has $w=-1$, and since $\Psi_{ABCD}$ has $w=0$, then $\psi_{AB}$ has $w=-1$. Thus:
@@
[…]
@@
so the term with $S$ in (3.6) vanishes. Suppose that, in addition, $(M,g_{ab})$ is Hermitian, with complex structure $J^{a}{}_{b}=\sqrt{2}Z_{bc}g^{ca}$ (so that the principal spinors are $o_{A},\iota_{A}$, recall (3.1) and (2.1c)). Then from appendix A, we have $\sigma^{0}_{A^{\prime}}=0$ and the term with $*K$ in (3.6) vanishes. Recalling that for an Einstein-Hermitian geometry with $\Psi_{2}\neq 0$, the Lee form is (2.15), the result (3.6) reduces to
@@
One way of interpreting this identity is in terms of spin-lowering via Killing spinors [40, §6.4]. We have $\Psi_{2}^{-1/3}\Sigma_{ab}\propto\Psi_{ABCD}K^{CD}\epsilon_{A^{\prime}B^{\prime}}$, where $K_{AB}=\Psi_{2}^{-1/3}o_{(A}\iota_{B)}$ is a valence-2 Killing spinor, $\nabla_{A^{\prime}(A}K_{BC)}=0$ (cf. [40, Eq. (6.7.15)]). Then (3.9) is the statement that
@@
is an anti-self-dual Maxwell field. As explained by Penrose and Rindler in [40, §6.4], this construction can be applied to linearized gravity on Minkowski space-time by replacing $\Psi_{ABCD}$ with a spin 2 zero-rest-mass field. Then one can associate a charge to this field by integrating (3.10) on a closed 2-surface, where, since Minkowski space-time is topologically trivial, the 2-surface must surround a region of matter sources in order for the charge to be non-vanishing.
@@ 3.2. Perturbations and closed 2-forms
We say that $T=T(s)\big{|}_{s=0}$ is the “background” value of $T(s)$, and $\delta{T}$ is the linearization of $T(s)$ at $T$. We recall that, under an infinitesimal diffeomorphism generated by an arbitrary vector field $\xi^{a}$, the linearization $\delta{T}$ transforms as
@@
The Levi-Civita connection of $g_{ab}(s)$ is denoted $\nabla^{(s)}_{a}$, and its relation to the background Levi-Civita connection is (see [41, Eq. (7.5.7)])
@@
Finally, consider a gauge transformation $\delta\hat{\omega}\to\delta\hat{\omega}+\pounds_{\xi}\hat{\omega}$. We have £$\pounds_{\xi}\hat{\omega}={\rm d}(\xi\lrcorner\,\hat{\omega})$ since $\hat{\omega}$ is closed. Hence, $\delta\hat{\omega}\to\delta\hat{\omega}+{\rm d}(\xi\lrcorner\,\hat{\omega})$, so the result (3.20) is gauge-invariant. ∎
@@ 4. Examples
A non-trivial issue is how to find the independent 2-cycles, for which we need the second homology group $H_{2}(M)$. To do this, we shall use the fact that all of the examples are toric. This implies that one can apply the rod structure formalism as in [15], which allows to determine the topological properties of the instanton. For example, for a rod structure with $n$ turning points, the Euler characteristic $\chi(M)$ is equal to $n$. In the non-compact case, we see from [36, Theorem 1] that $H_{2}(M)\cong\mathbb{Z}^{n-1}$. In the compact case, we can use the fact that $\chi(M)$ is also the alternating sum of the Betti numbers $b_{i}(M)$ to deduce the number of independent 2-cycles.
Below we shall also use the terminology of ‘nuts’ and ‘bolts’, for which we refer to the paper by Gibbons and Hawking [24].
@@ 4.1. Ambi-Kähler examples
define integrable almost-complex structures with opposite orientation. The Kähler forms are then given by $\hat{\omega}^{\pm}=(\Psi_{2}^{\pm})^{2/3}\omega^{\pm}$. The topology is $\mathbb{R}^{2}\times S^{2}$, and the rod structure has two turning points, two semi-infinite rods and one finite rod [15, §4.4]. Since there are two turning points, we have $H_{2}(M)\cong\mathbb{Z}$ [36, Thm. 1], so there is only one independent 2-cycle $S$, corresponding to the bolt at $r=r_{+}$ of the Killing field $\partial_{\tau}+\Omega_{E}\partial_{\phi}$, where $r_{+}=m+\sqrt{m^{2}+a^{2}}$ and $\Omega_{E}=a/(2mr_{+})$. The associated charges (1.3) are
[…]
@@
The self-dual Taub-NUT instanton is obtained from (4.5) when $m=n$. It is ALF, and the topology is $S^{4}\setminus\{\rm pt.\}\cong\mathbb{R}^{4}$. From (4.6), we see that $\Psi^{-}_{2}=0$ and $\Psi^{+}_{2}=2n/(r+n)^{3}$, that is $W^{-}_{abcd}=0$ and $W^{+}_{abcd}\neq 0$, hence the space is hyper-Kähler with respect to one orientation, i.e. there is a 2-sphere of Kähler structures, and type D with respect to the other. The 2-forms (4.7) are however not Kähler; the geometry is not only hyper-Kähler but also strictly conformally Kähler w.r.t. both orientations [7]. The rod structure has one turning point and two semi-infinite rods [15, §4.2]. However, since the manifold is topologically trivial, there are no 2-cycles; thus the parameter $n$ does not correspond to a charge integral like (1.3). For a different construction involving 3-cycles, see [13].
@@
The Taub-bolt instanton is obtained from (4.5) when $m=\frac{5}{4}n$. It is ALF, and the topology is $\mathbb{CP}^{2}\setminus\{{\rm pt.}\}$. We have $\Psi_{2}^{-}=n/[4(r-n)^{3}]$, $\Psi_{2}^{+}=9n/[4(r+n)^{3}]$, so both $W^{\pm}_{abcd}$ remain type D. The rod structure has two turning points, two semi-infinite rods and one finite rod [15, §4.7]. As in Kerr we have two turning points, so $H_{2}(M)\cong\mathbb{Z}$ [36, Thm. 1] and there is one independent 2-cycle, which we denote by $S_{2n}$ and corresponds to the bolt of $\partial_{\psi}$ at $r=2n$. We then find:
@@
where $f(r)=1-(a/r)^{4}$, and $a$ is a real constant. The topology is $T^{*}\mathbb{CP}^{1}$, and the asymptotics is ALE. The metric is Ricci-flat, the self-dual Weyl tensor is $W^{+}_{abcd}=0$, and the anti-self-dual Weyl tensor $W^{-}_{abcd}$ is type D  [^4] As noted by Gibbons and Pope [27], the Eguchi-Hanson metric (4.9) can be obtained from (4.5) by setting $m=n+a^{4}/(128n^{3})$, $r=m+\tilde{r}^{2}/(8n)$, and taking the limit $n\to\infty$ (and then redefining $\tilde{r}\to r$), see [27, Eqs. (4.15)-(4.16)]. This procedure leads to (4.10) (i.e., $W^{-}_{abcd}\neq 0$ and $W^{+}_{abcd}=0$, which is the opposite to self-dual Taub-NUT).:
@@
give integrable almost-complex structures $J^{\pm}$ with opposite orientation, and that ${\rm d}\omega^{+}=0$, so $J^{+}$ is Kähler. However ${\rm d}\omega^{-}\neq 0$, and this gives the conformally Kähler form $\hat{\omega}^{-}=(\Psi_{2}^{-})^{2/3}\omega^{-}$ needed to define the charge (1.3). The rod structure has two turning points, two semi-infinite rods and one finite rod [15, §4.5]. As in previous cases, two turning points give $H_{2}(M)\cong\mathbb{Z}$ [36, Thm. 1] so there is one independent 2-cycle, which we denote by $S_{a}$ and corresponds to the bolt of $\partial_{\psi}$ at $r=a$. We have
@@ 4.2. The Chen-Teo instanton
The Chen-Teo family of ALF Ricci-flat metrics is [17, Eq. (2.1)]
@@
The scalar $\Psi_{2}$ can be computed using [8, Eq. (2.28)]; we find
@@
The vector fields $\partial_{\tau},\partial_{\phi}$ are Killing, and the solution is toric. The rod structure has three turning points, two finite rods and two semi-infinite rods [17, 2, 31]. These can be conveniently described in terms of the four roots, $r_{1},r_{2},r_{3},r_{4}$, of the polynomial $P$ in (4.19). Following Chen and Teo [17, §3.3], we have
@@
where the notation $\ell_{i}=(\alpha_{i},\beta_{i})$ means $\ell_{i}=\alpha_{i}\partial_{\tau}+\beta_{i}\partial_{\phi}$, and the expressions for $\alpha_{i},\beta_{i}$ can be found in [17, Eqs. (3.14)-(3.15)]. It is convenient to adapt the Killing coordinates to the rod vectors $\ell_{1},\ell_{2}$ (cf. [2]). This is done by defining $(\tilde{\tau},\tilde{\phi})$ via $\tau=\alpha_{2}\tilde{\tau}+\alpha_{1}\tilde{\phi}$, $\phi=\beta_{2}\tilde{\tau}+\beta_{1}\tilde{\phi}$, so that $\partial_{\tilde{\phi}}=\ell_{1}$ and $\partial_{\tilde{\tau}}=\ell_{2}$. The closed Kähler form $\hat{\omega}=\Psi_{2}^{2/3}\omega$ is then
@@
where we used the notation $\alpha_{i}=K_{i}[1]/\kappa_{i}$ and $\beta_{i}=1/\kappa_{i}$ following [17, Eqs. (3.14)-(3.15)].
The family of ALF metrics (4.18) has, in general, conical singularities. Chen and Teo showed in [17, §4.3] that, by imposing certain constraints on the parameters, one can obtain a smooth, AF, regular 4-manifold. This is the Chen-Teo instanton [16], which corresponds to setting the total NUT charge to zero and requiring the rod vectors to have closed orbits with $2\pi$ period and to satisfy $\ell_{1}=\ell_{4}=\pm\ell_{2}\pm\ell_{3}$, cf. [17, Eq. (4.13)] and [2]. The solution to these constraints is given in [17, Eq. (4.16)], from where one deduces that
@@
The resulting family of metrics has 2 parameters, $(\xi,k)$. The parametrization (4.23) gives $\alpha_{1}=\alpha_{4}$, $\beta_{1}=\beta_{4}$, $\alpha_{3}=-(\alpha_{1}+\alpha_{2})$ $\beta_{3}=-(\beta_{1}+\beta_{2})$. The rod vectors are then $\ell_{1}=\ell_{4}=\partial_{\tilde{\phi}}$, $\ell_{2}=\partial_{\tilde{\tau}}$, $\ell_{3}=-(\partial_{\tilde{\phi}}+\partial_{\tilde{\tau}})$.
The topology is $\mathbb{CP}^{2}\setminus S^{1}$. Since there are three turning points, we have $H_{2}(M)\cong\mathbb{Z}^{2}$ [36, Thm. 1], so there are two independent 2-cycles, which we can take to be the finite rods $\mathcal{R}_{2},\mathcal{R}_{3}$. Using the expression (4.22), it is straightforward to integrate $\hat{\omega}$ over $\mathcal{R}_{2},\mathcal{R}_{3}$. Since the rods correspond to roots of $X,Y$, the terms with $H$ in (4.22) vanish on the rods, and also (from (4.19)) $G/F$ becomes constant, so we find (using $a_{4}=0$, $a_{3}=1$)
@@
This result implies that both of the two parameters of the Chen-Teo instanton can be obtained as charge integrals of the Kähler form. As mentioned in the introduction, this is in contrast to the Kerr case, where the Kähler construction gives only one of the two parameters, namely the mass, see (4.4). We can understand this difference from the fact that there are two bolts in Chen-Teo, whereas Kerr has only one. In a sense, the Chen-Teo instanton is a “double” object, as suggested by the interpretation in [17, §5] in terms of two touching Kerr-NUTs.
5. Linearized Kähler structures
A compact Einstein metric with the property $\det(W^{+})>0$ is necessarily conformally Kähler, a remarkable result shown by Wu in [42]. LeBrun’s alternative proof [34] can be adapted to manifolds with a different asymptotic structure as in [12, Proposition 3]. This in turn was used in [12] to show that Ricci-flat metrics sufficiently close to a Ricci-flat, ALF, toric, Hermitian instanton must also be Hermitian.
In this section we shall use ideas from [34, 12] to deduce some identities that allow us to interpret our basic result (3.20) as the linearized version of a complex structure for the perturbed instanton.

### v3

@@ 2.2. The self-dual Weyl spinor
In section 5 we shall be interested in eigenspinors for $\Psi_{ABCD}$. A detailed study of this can be found in [51, §8.3]. Here we collect some facts adapted to Riemannian signature, for later use in section 5.
@@
where the $\phi^{i}_{AB}'$s are orthonormal, $I^{ABCD}\phi^{i}_{AB}\phi^{j}_{CD}=\delta^{ij}$. From [51, Eq. (8.3.8), (8.3.11)], the eigenvalues satisfy the following identities:
@@ 3.1. Charges for Hermitian instantons
The proof of (3.3) is straightforward if one uses a `conformal-GHP connection' [7]: this is an operator $\mathcal{C}_{AA^{\prime}}$ which generalizes the ordinary GHP connection [50, §4.12] to be invariant under combined conformal transformations and rescalings of the spinors $o_{A},\iota_{A}$, by taking into account both the conformal and the GHP weights of the objects it acts on. The necessary background is given in appendix A. Since $\zeta_{AB}$ has conformal weight $w=1$, then $\zeta^{AB}$ has w=1w=-1, and since $\Psi_{ABCD}$ has $w=0$, then $\psi_{AB}$ has w=1w=-1. Thus:
@@
$={}$ $(\mathcal{C}_{A^{\prime}}{}^{B}\Psi_{ABCD})\zeta^{CD}+\Psi_{ABCD}\mathcal{C}_{A^{\prime}}{}^{B}\zeta^{CD}$
@@
$={}$ $f_{A^{\prime}}{}^{B}\psi_{AB}+\zeta^{CD}\nabla_{A^{\prime}}{}^{B}\Psi_{ABCD}+K_{AA^{\prime}},$
@@
so the term with $S$ in (3.6) vanishes. Suppose that, in addition, $(M,g_{ab})$ is Hermitian, with complex structure $J^{a}{}_{b}=\sqrt{2}Z_{bc}g^{ca}$ (so that the principal spinors are $o_{A},\iota_{A}$, recall (3.1) and (2.1c)). Then from appendix A, we have $\sigma^{0}_{A^{\prime}}=0$ and the term with K*K in (3.6) vanishes. Recalling that for an Einstein-Hermitian geometry with $\Psi_{2}\neq 0$, the Lee form is (2.15), the result (3.6) reduces to
@@
One way of interpreting this identity is in terms of spin-lowering via Killing spinors [51, §6.4]. We have $\Psi_{2}^{-1/3}\Sigma_{ab}\propto\Psi_{ABCD}K^{CD}\epsilon_{A^{\prime}B^{\prime}}$, where $K_{AB}=\Psi_{2}^{-1/3}o_{(A}\iota_{B)}$ is a valence-2 Killing spinor, $\nabla_{A^{\prime}(A}K_{BC)}=0$ (cf. [51, Eq. (6.7.15)]). Then (3.9) is the statement that
@@
is an anti-self-dual Maxwell field. As explained by Penrose and Rindler in [51, §6.4], this construction can be applied to linearized gravity on Minkowski space-time by replacing $\Psi_{ABCD}$ with a spin 2 zero-rest-mass field. Then one can associate a charge to this field by integrating (3.10) on a closed 2-surface, where, since Minkowski space-time is topologically trivial, the 2-surface must surround a region of matter sources in order for the charge to be non-vanishing.
@@ 3.2. Perturbations and closed 2-forms
We say that $T=T(s)\big|_{s=0}$ is the ``background'' value of $T(s)$, and $\delta{T}$ is the linearization of $T(s)$ at $T$. We recall that, under an infinitesimal diffeomorphism generated by an arbitrary vector field $\xi^{a}$, the linearization $\delta{T}$ transforms as
@@
The Levi-Civita connection of $g_{ab}(s)$ is denoted $\nabla^{(s)}_{a}$, and its relation to the background Levi-Civita connection is (see [54, Eq. (7.5.7)])
@@
[…]
@@ 4. Examples
A non-trivial issue is how to find the independent 2-cycles, for which we need the second homology group $H_{2}(M)$. To do this, we shall use the fact that all of the examples are toric. This implies that one can apply the rod structure formalism as in [19], which allows to determine the topological properties of the instanton. For example, for a rod structure with $n$ turning points, the Euler characteristic $\chi(M)$ is equal to $n$. In the non-compact case, we see from [47, Theorem 1] that $H_{2}(M)\cong\mathbb{Z}^{n-1}$. In the compact case, we can use the fact that $\chi(M)$ is also the alternating sum of the Betti numbers $b_{i}(M)$ to deduce the number of independent 2-cycles.
Below we shall also use the terminology of `nuts' and `bolts', for which we refer to the paper by Gibbons and Hawking [32].
@@ 4.1. Ambi-Kähler examples
define integrable almost-complex structures with opposite orientation. The Kähler forms are then given by $\hat{\omega}^{\pm}=(\Psi_{2}^{\pm})^{2/3}\omega^{\pm}$. The topology is $\mathbb{R}^{2}\times S^{2}$, and the rod structure has two turning points, two semi-infinite rods and one finite rod [19, §4.4]. Since there are two turning points, we have $H_{2}(M)\cong\mathbb{Z}$ [47, Thm. 1], so there is only one independent 2-cycle $S$, corresponding to the bolt at $r=r_{+}$ of the Killing field $\partial_{\tau}+\Omega_{E}\partial_{\phi}$, where $r_{+}=m+\sqrt{m^{2}+a^{2}}$ and $\Omega_{E}=a/(2mr_{+})$. The associated charges (1.4) are
$\int_{S}\hat{\omega}^{\pm}=\pm\,4\pi\,m^{2/3}.$ (4.4)
@@
The self-dual Taub-NUT instanton is obtained from (4.5) when $m=n$. It is ALF, and the topology is $S^{4}\setminus\{\rm pt.\}\cong\mathbb{R}^{4}$. From (4.6), we see that $\Psi^{-}_{2}=0$ and $\Psi^{+}_{2}=2n/(r+n)^{3}$, that is $W^{-}_{abcd}=0$ and $W^{+}_{abcd}\neq 0$, hence the space is hyper-Kähler with respect to one orientation, i.e. there is a 2-sphere of Kähler structures, and type D with respect to the other. The 2-forms (4.7) are however not Kähler; the geometry is not only hyper-Kähler but also strictly conformally Kähler w.r.t. both orientations [8]. The rod structure has one turning point and two semi-infinite rods [19, §4.2]. However, since the manifold is topologically trivial, there are no 2-cycles; thus the parameter $n$ does not correspond to a charge integral like (1.4). For a different construction involving 3-cycles, see [17].
@@
The Taub-bolt instanton is obtained from (4.5) when $m=\frac{5}{4}n$. It is ALF, and the topology is $\mathbb{CP}^{2}\setminus\{{\rm pt.}\}$. We have $\Psi_{2}^{-}=n/[4(r-n)^{3}]$, $\Psi_{2}^{+}=9n/[4(r+n)^{3}]$, so both $W^{\pm}_{abcd}$ remain type D. The rod structure has two turning points, two semi-infinite rods and one finite rod [19, §4.7]. As in Kerr we have two turning points, so $H_{2}(M)\cong\mathbb{Z}$ [47, Thm. 1] and there is one independent 2-cycle, which we denote by $S_{2n}$ and corresponds to the bolt of $\partial_{\psi}$ at $r=2n$. We then find:
@@
where $f(r)=1-(a/r)^{4}$, and $a$ is a real constant. The topology is $T^{*}\mathbb{CP}^{1}$, and the asymptotics is ALE. The metric is Ricci-flat, the self-dual Weyl tensor is $W^{+}_{abcd}=0$, and the anti-self-dual Weyl tensor $W^{-}_{abcd}$ is type D  [^4] As noted by Gibbons and Pope [35], the Eguchi-Hanson metric (4.9) can be obtained from (4.5) by setting $m=n+a^{4}/(128n^{3})$, $r=m+\tilde{r}^{2}/(8n)$, and taking the limit $n\to\infty$ (and then redefining $\tilde{r}\to r$), see [35, Eqs. (4.15)-(4.16)]. This procedure leads to (4.10) (i.e., $W^{-}_{abcd}\neq 0$ and $W^{+}_{abcd}=0$, which is the opposite to self-dual Taub-NUT).:
@@
give integrable almost-complex structures $J^{\pm}$ with opposite orientation, and that ${\rm d}\omega^{+}=0$, so $J^{+}$ is Kähler. However ${\rm d}\omega^{-}\neq 0$, and this gives the conformally Kähler form $\hat{\omega}^{-}=(\Psi_{2}^{-})^{2/3}\omega^{-}$ needed to define the charge (1.4). The rod structure has two turning points, two semi-infinite rods and one finite rod [19, §4.5]. As in previous cases, two turning points give $H_{2}(M)\cong\mathbb{Z}$ [47, Thm. 1] so there is one independent 2-cycle, which we denote by $S_{a}$ and corresponds to the bolt of $\partial_{\psi}$ at $r=a$. We have
@@ 4.2. The Chen-Teo instanton
The Chen-Teo family of ALF Ricci-flat metrics is [21, Eq. (2.1)]
@@
The scalar $\Psi_{2}$ can be computed using [9, Eq. (2.28)]; we find
@@
The vector fields $\partial_{\tau},\partial_{\phi}$ are Killing, and the solution is toric. The rod structure has three turning points, two finite rods and two semi-infinite rods [21, 2, 41]. These can be conveniently described in terms of the four roots, $r_{1},r_{2},r_{3},r_{4}$, of the polynomial $P$ in (4.19). Following Chen and Teo [21, §3.3], we have
@@
[…]
@@
where we used the notation $\alpha_{i}=K_{i}[1]/\kappa_{i}$ and $\beta_{i}=1/\kappa_{i}$ following [21, Eqs. (3.14)-(3.15)].
The family of ALF metrics (4.18) has, in general, conical singularities. Chen and Teo showed in [21, §4.3] that, by imposing certain constraints on the parameters, one can obtain a smooth, AF, regular 4-manifold. This is the Chen-Teo instanton [20], which corresponds to setting the total NUT charge to zero and requiring the rod vectors to have closed orbits with $2\pi$ period and to satisfy $\ell_{1}=\ell_{4}=\pm\ell_{2}\pm\ell_{3}$, cf. [21, Eq. (4.13)] and [2]. The solution to these constraints is given in [21, Eq. (4.16)], from where one deduces that
@@
[…]
The topology is $\mathbb{CP}^{2}\setminus S^{1}$. Since there are three turning points, we have $H_{2}(M)\cong\mathbb{Z}^{2}$ [47, Thm. 1], so there are two independent 2-cycles, which we can take to be the finite rods $\mathcal{R}_{2},\mathcal{R}_{3}$. Using the expression (4.22), it is straightforward to integrate $\hat{\omega}$ over $\mathcal{R}_{2},\mathcal{R}_{3}$. Since the rods correspond to roots of $X,Y$, the terms with $H$ in (4.22) vanish on the rods, and also (from (4.19)) $G/F$ becomes constant, so we find (using $a_{4}=0$, $a_{3}=1$)
@@
This result implies that both of the two parameters of the Chen-Teo instanton can be obtained as charge integrals of the Kähler form. As mentioned in the introduction, this is in contrast to the Kerr case, where the Kähler construction gives only one of the two parameters, namely the mass, see (4.4). We can understand this difference from the fact that there are two bolts in Chen-Teo, whereas Kerr has only one. In a sense, the Chen-Teo instanton is a ``double'' object, as suggested by the interpretation in [21, §5] in terms of two touching Kerr-NUTs.
5. Infinitesimal rigidity
In this section we shall use ideas from [44, 16] to deduce some identities that allow us to interpret our basic result (3.20) as the linearized version of a complex structure for the perturbed instanton. These will be used to analyze the moduli space of Einstein metrics and prove our main theorem 1.1.
Recall that, given a smooth manifold $M$, the moduli space $\mathscr{E}(M)$ is the quotient of the set of Einstein metrics on $M$ by the diffeomorphism group ${\rm Diff}(M)$. The space $\mathscr{E}(M)$ is not necessarily a manifold, as it might contain singular points that make it a non-smooth variety. One is thus interested in achieving a local description of $\mathscr{E}(M)$ around some fixed Einstein metric $g\in\mathscr{E}(M)$. If $g$ has no infinitesimal Einstein deformations, it is said to be rigid or isolated in $\mathscr{E}(M)$. If $g$ does have infinitesimal deformations, a natural question arises: are these deformations tangent to a curve of Einstein metrics in $\mathscr{E}(M)$? In other words, does any infinitesimal deformation integrate to a genuine Einstein metric? If this is the case, then $\mathscr{E}(M)$ is said to be integrable at $g$.
It is a notable open problem whether integrability holds in general for Einstein metrics, cf. [12, 13]. Examples where integrability holds include special, compact geometries such as Calabi-Yau and Kähler-Einstein [12, Chapter 12], and moduli spaces of self-dual Yang-Mills connections on self-dual compact manifolds [11]. On the other hand, the canonical example of non-integrability was given by Koiso [40], who showed that the standard metric on $M=\mathbb{CP}^{1}\times\mathbb{CP}^{2n}$ has infinitesimal deformations that do not integrate to curves in $\mathscr{E}(M)$. Other examples arise for instance in the Lorentzian case, where integrability in the above sense is referred to as `linearization stability' [27, 28]: Fischer, Marsden and Moncrief [29] showed that, if $(M,g)$ is a vacuum space-time with a compact Cauchy surface and a Killing field, then $\mathscr{E}(M)$ has the structure of a quadratic variety near $g$ —i.e., $\mathscr{E}(M)$ has a conical singularity at $g$.
In our analysis, we shall use ideas due to Wu [55] and LeBrun [44]. Wu proved the remarkable result that a compact Einstein metric with the property $\det(W^{+})>0$ is necessarily conformally Kähler. An alternative proof was given by LeBrun, that can be adapted to manifolds with a different asymptotic structure as in [16, Proposition 3]. This in turn was used in [16] to show that Ricci-flat metrics sufficiently close to a Ricci-flat, ALF, toric, Hermitian instanton must also be Hermitian.